# thanks: Corresponding author

### v2

@@ II.1.2 Parametrization of the distribution $q=(q_{0},\ldots,q_{d-1})$
For an arbitrary $d$-outcome probability distribution $q=(q_{0},\ldots,q_{d-1})$, we define a (d1)(d-1)-dimensional Hermitian matrix $N$ such that it has nonzero entries only along the antidiagonal, i.e., $N_{kk^{\prime}}\neq 0$ if and only if $k\oplus k^{\prime}=d$. More precisely, let $N_{kk^{\prime}}=\sum_{m=0}^{d-1}\omega^{k^{\prime}m}q_{m}=\sum_{m=0}^{d-1}\omega^{-km}q_{m}$, so that for nonzero entries we have
@@
The matrix $N$ captures the non-uniformity of the probability distribution $q=(q_{0},\ldots,q_{d-1})$, in the sense that $N=0$ if and only if the underlying distribution is uniform, i.e., $q=(1/d,\ldots,1/d)$. Clearly, $N$ has $\lceil\frac{d-1}{2}\rceil$ independent entries, where $\lceil\frac{d-1}{2}\rceil$ denotes the smallest integer greater than or equal to $\frac{d-1}{2}$; that is, it equals $\frac{d-1}{2}$ if $d$ is odd, and $\frac{d}{2}$ if $d$ is even. All entries are complex, except for $N_{d/2,d/2}$ which is real and occurs when $d$ is even. With this notation, the (d1)(d-1)-dimensional matrix $N$ is fully determined by d1d-1 real independent parameters. Accordingly, by writing $N_{k,d-k}=\delta_{2k-1}+i\delta_{2k}$, we have for the (d1)(d-1) real parameters $\delta_{1},\ldots,\delta_{d-1}$
@@
for $k=1,\ldots,\lceil\frac{d-1}{2}\rceil$. This parametrization is particularly useful to quantify the extent to which the probability distribution $q=(q_{0},\ldots,q_{d-1})$ deviates from the uniform distribution $q=(1/d,\ldots,1/d)$. To illustrate the nature of this parametrization, we explicitly consider two cases of $d=2$ and $d=3$.
For $d=2$, the only nonzero entry of $N$ is given by $N_{11}=\delta$, where $\delta=q_{0}-q_{1}$, such that $1\leq\delta\leq 1$. It follows therefore that
@@
II.2 Mutually equi-biased bases for $d=2$
For the simplest case of two-level system, suppose $\mathcal{B}^{(1)}=\{|0\rangle,|1\rangle\}$ represents the standard basis. Associated with the probability distribution ${q}=(q_{0},q_{1})$, or its corresponding bistochastic matrix $\mathcal{Q}=\begin{pmatrix}q_{0}&q_{1}\\ q_{1}&q_{0}\end{pmatrix}$, we have to construct the unitary matrix $\mathcal{U}$ with prescribed moduli of its matrix elements. Starting with $\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}\\ \sqrt{q_{1}}&\sqrt{q_{0}}{\mathrm{e}}^{i\xi}\end{pmatrix}$, unitarity requires $\xi=\pi$, hence
@@ II.2 Mutually equi-biased bases for $d=2$
Although both $\mathcal{U}$ and $\mathcal{V}$ generate the same bistochastic matrix $\mathcal{Q}$ and both define bases that are MEBs with respect to the standard basis, they are not necessarily MEBs with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equi-biased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same bistochastic matrix $\mathcal{Q}$. This places a constraint on $\alpha$ such that $q_{0}^{2}+q_{1}^{2}+2q_{0}q_{1}\cos{\alpha}=q_{0}$, where uniquely determines $\alpha$ as
@@
where the unbiasedness parameter $\delta=q_{0}-q_{1}$ is defined by Eq. (7).
Starting from the standard orthonormal basis $\{|e_{k}^{(1)}\rangle=|k\rangle\}$, one can apply the unitary matrices $\mathcal{U}$ and $\mathcal{V}$ to construct new bases $\{|e_{k}^{(2)}\rangle=\mathcal{U}|e_{k}^{(1)}\rangle\}$ and $\{|e_{k}^{(3)}\rangle=\mathcal{V}|e_{k}^{(1)}\rangle\}$, for $k=0,1$. This notion of equi-biasedness admits an interesting geometric interpretation in the Bloch vector representation, in the sense that with $P_{k}^{(\alpha)}=({\mathbb{I}}_{2}+\boldsymbol{r}_{k}^{(\alpha)}\cdot\boldsymbol{\sigma})/2$ as the Bloch vector representation of the projection operator $P_{k}^{(\alpha)}$, the Bloch vectors satisfy
@@
Figure 1: Bloch vector representation for MEB, when $d=2$. Geometrically, equi-biasedness means equal angle, $\vartheta=\cos^{-1}[\delta]$, between Bloch vectors of different bases.
II.3 Mutually equi-biased bases for $d=3$
@@
The unitarity condition $\mathcal{U}\mathcal{U}^{\dagger}={\mathbb{I}}$, i.e., the mutual orthogonality of the rows (and columns), gives simply the following conditions
@@
Equations (19) to (21) imply that the unitarity is fulfilled if we find the phases in such a way that three line segments $L_{0},L_{1},L_{2}$ form a triangle. This imposes the triangle inequality on the line segments ZyczkowskiCMP2005
@@
where $(i,j,k)$ are different choices of $(0,1,2)$. For probabilities $q=(q_{0},q_{1},q_{2})$ for which the above inequalities hold, Eqs. (19) to (21) give the phases as
@@
For each probability distribution $q=(q_{0},q_{1},q_{2})$ satisfying the chain-link conditions (23), the associated unitary matrix $\mathcal{U}$, defined in Eq. (18) with phases specified in Eqs. (24) to (26), generates the corresponding unistochastic matrix $\mathcal{Q}$. By construction, the matrix $\mathcal{U}$ represents the most general form consistent with $q=(q_{0},q_{1},q_{2})$ and defines the second orthonormal basis $\{|e_{k}^{(2)}\rangle=\mathcal{U}|e_{k}^{(1)}\rangle\}$ with respect to the standard basis $\{|e_{k}^{(1)}\rangle\}$. Figure 2(a) shows probability triangle with its unistochastic subset. The boundaries of this subset are hypocycloids, each corresponding to the saturation of one of the chain-link inequalities (23). A complete clarification of Fig. 2, along with the additional conditions leading to the feasible region, is provided in Remark 3.
Given $\mathcal{U}$, we have to construct two unitary matrices, $\mathcal{V}$ and $\mathcal{W}$, such that all bases are mutually equi-biased. The following proposition provides a complete set of four MEBs for $d=3$.
Figure 2: Geometric representation of the probability distribution $q=(q_{0},q_{1},q_{2})$ and the construction of the feasible region. (a) Equilateral triangle associated to probability distribution $q=(q_{0},q_{1},q_{2})$, with its unistochastic subset. All subfigures are plotted in the $\delta_{2}\delta_{1}$-plane. Each subfigure corresponds to a constraint applied in succession. See Remark 3 for details.
@@
Starting from the standard basis $\{|e_{k}^{(1)}\rangle\}_{k=0}^{2}$, the unitary matrices $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{W}$ generate a complete set of four MEBs. Here, $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ and $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$, where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$ and $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$, respectively, and $\mathcal{U}$ is defined by Eq. (18). The pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$ are distinct and take their values from the following set
@@
Before proceeding with the proof, we illustrate the form of the solutions in Fig. 3. They represent a one-parameter class of probability distributions $q=(q_{0},q_{1},q_{2})$ within the feasible region of the probability triangle, defined by six curves each starting at $\mu=1/3$ at the center of the feasible region and ending at $\mu=1/2$ on the boundary. The curves represent the allowed values of the distribution $q=(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$, or any permutation of its probabilities, for $(\alpha_{1},\alpha_{2})$ chosen from the pairs given in Eq. (27).
@@
Figure 3: Geometrical representation of the probability distribution $q=(q_{0},q_{1},q_{2})$, for which a complete set of four MEBs exist. The figure is plotted in the $\delta_{2}\delta_{1}$-plane, where independent parameters $\delta_{1}$ and $\delta_{2}$ are defined in Eq. (8). The curves represent the allowed values of the distribution $q=(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$, or any permutation of its probabilities, when $\alpha_{1}=0$ and $\alpha_{2}=\cos^{-1}[-\mu/(1-\mu)]$. The circles correspond to regions with constant index $\mu$. Corresponding to each $\mu$, only six points, i.e., the distribution $q$ and its permutations, are allowed for a given $\mu$.
Proof.
Having the first unitary matrix $\mathcal{U}$, we have to find two other unitary matrices, $\mathcal{V}$ and $\mathcal{W}$, with the same moduli of matrix elements as $\mathcal{U}$. The second unitary matrix $\mathcal{V}$ can be constructed as $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$. Regarding the probability $q=(q_{0},q_{1},q_{2})$, although both $\mathcal{U}$ and $\mathcal{V}$ generate bases that are equi-biased with respect to the standard basis, they are not necessarily equi-biased with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equi-biased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same probability matrix $\mathcal{Q}$, up to a permutation of its entries. This places some constraints on phases $\alpha_{1}$ and $\alpha_{2}$.
Here, $\sigma\in S_{3}$ denotes a permutation of three elements $\{0,1,2\}$, where $S_{3}$ is the symmetric group of degree $3$. For clarity, the six operations of $S_{3}$ can be written as: the identity $\mathrm{id}$, the transpositions $(01)$, $(02)$, $(12)$, and the forward and backward cycles $(012)$ and $(210)$. After some calculations, we get
which holds for $\mu\in[1/3,5/9]$.
Equation (35) imposes a constraint on the allowed values of the phases $\alpha_{1}$ and $\alpha_{2}$ under which two unitary matrices $\mathcal{U}$ and $\mathcal{V}$ are mutually equi-biased. A contour plot of this equation in $\alpha_{1}\alpha_{2}$-plane for different values of $\mu$ is illustrated in Fig. 4. The contours which are bounded within the hexagon correspond to $\mu\in[1/3,1/2]$, and those outside the hexagon, i.e., within the two outer triangles, correspond to $\mu\in(1/2,5/9]$.
Figure 4: Contour plot of the lhs of Eq. (35) in $\alpha_{1}\alpha_{2}$-plane for $\alpha_{1},\alpha_{2}\in(-\pi,+\pi)$. (a) The inner boundary (green-ellipse) corresponds to $\mu=1/3$, (b) the outer boundary (red-hexagon) corresponds to $\mu=1/2$, and (c) two little pink contours are plotted for $\mu\approx 5/9$. The exact value for $\mu=5/9$ correspond to two points at $(2\pi/3,-2\pi/3)$ and $(-2\pi/3,2\pi/3)$. The $\alpha_{1}\alpha_{2}$-plane is divided by dashed grid lines to make it easier to identify the quadrant of the angles $\alpha_{1}$ and $\alpha_{2}$.
Clearly, the constraint imposed by Eq. (35) is invariant under the symmetric group $S_{3}$; it depends only on $\mu$ but not on where the probability $q=(q_{0},q_{1},q_{2})$ is located in the feasible region. However, a look at the LHSs of Eqs. (32) to (34) reveals that this set is invariant under a subgroup of $S_{3}$, namely, the cyclic group $C_{3}=\{\mathrm{id},(012),(210)\}$. The set, however, will possess the full symmetry under $S_{3}$ if and only if two out of the three terms $\cos{\alpha_{1}}$, $\cos{\alpha_{2}}$, and $\cos{\alpha_{12}}$ become equal. This results in three possible cases as
  1. (a) :
    $\alpha_{1}=0,\quad\quad\alpha_{2}=\theta^{\mu}$,
  2. (b) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=0$,
  3. (c) :
    $\alpha_{1}=\theta^{\mu},\quad\;\alpha_{2}=\theta^{\mu}$,
where, from Eq. (35), $\theta^{\mu}$ is obtained as Eq. (28). The index condition $\mu\leq 1/2$ is obtained from $|\cos{\theta^{\mu}}|\leq 1$. Using this in Eqs. (32) to (34), we get for the case (a) the following set of equations
Depending on the choice of the permutation $\sigma\in S_{3}$, the set of equations determines the probability distribution $q=(q_{0},q_{1},q_{2})$. More precisely, the choice $\sigma=\mathrm{id}$ fixes $q_{1}=q^{\mu}_{0}$, so that $q=(q^{\mu}_{\pm},q^{\mu}_{0},q^{\mu}_{\mp})$, where $q^{\mu}_{0}=(1-\mu)/2$ and $q^{\mu}_{\pm}$ is given in Eq. (30). Similarly, the cycle $\sigma=(012)$ fixes $q_{0}=q^{\mu}_{0}$, giving $q=(q^{\mu}_{0},q^{\mu}_{\pm},q^{\mu}_{\mp})$, while the cycle $\sigma=(210)$ yields $q=(q^{\mu}_{\pm},q^{\mu}_{\mp},q^{\mu}_{0})$. What about the transpositions $(01)$, $(02)$, and $(12)$? In these cases, the set possesses solution only for $\mu=1/3$, corresponding to the uniform distribution $q=(1/3,1/3,1/3)$.
In a similar manner, for cases (b) and (c), the set of Eqs. (32) to (34) reduce to
and
respectively. In each case, the corresponding set of equations gives, up to a permutation, the same set of solutions (29) and (30) for the probabilities $(q_{0},q_{1},q_{2})$. Irrespective of which case (a), (b), or (c) is considered, only three out of six possible permutations $\sigma$, namely, those corresponding to the cyclic group $C_{3}=\{\mathrm{id},(012),(210)\}$, are contributed in the solutions. Table 1 presents all possible solutions of Eqs. (32) to (34), corresponding to three cases (a), (b), (c), and three permutations $\mathrm{id}$, $(012)$, $(210)$.
Table 1: Possible solutions of Eqs. (32) to (34).
$\sigma=\mathrm{id}$ $\sigma=(012)$ $\sigma=(210)$
To complete the characterization of the set of MEBs, we have to construct the third unitary matrix $\mathcal{W}$ as $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$ where $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$. Similar to the arguments given above, we require that the unitaries $\mathcal{W}\mathcal{U}$ and $\mathcal{W}\mathcal{V}$ generate the unistochastic matrix $\mathcal{Q}$, up to a permutation of its rows and columns. The first requirement is achieved if we select the pair $(\beta_{1},\beta_{2})$ from the cases (a), (b), or (c), mentioned above for $(\alpha_{1},\alpha_{2})$. To ensure that $\mathcal{W}\mathcal{V}$ also generates the unistochastic matrix $\mathcal{Q}$, we have to choose the pairs $(\beta_{1},\beta_{2})$ distinct from the pairs $(\alpha_{1},\alpha_{2})$. ∎
@@
It is clear from the construction that, although all three unitary matrices $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{W}$ generate the same unistochastic matrix $Q$, associated with the probability distribution $q=(q_{0},q_{1},q_{2})$, the unistochastic matrices resulting from their mutual products $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ are not necessary equal to $Q$. However, they are equivalent in the sense that they are generated either by $q=(q_{0},q_{1},q_{2})$ or by a permutation of its components, i.e., $q^{\sigma}=(q_{\sigma(0)},q_{\sigma(1)},q_{\sigma(2)})$, such as $q^{(012)}=(q_{1},q_{2},q_{0})$ and $q^{(210)}=(q_{2},q_{1},q_{0})$. All such possibilities are summarized in Table 2. The first column represents all possible choices for the distribution $q=(q_{0},q_{1},q_{2})$, and the first row represents all possible choices for two pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$. For each chosen probability and each chosen pairs of phases, three equivalent distributions for which three unitary matrices $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ generate equivalent unistochastic matrices are provided.
As we have seen, the mutual equi-biasedness imposes some constraints on the index $\mu=q_{0}^{2}+q_{1}^{2}+q_{2}^{2}$ of the distribution $q=(q_{0},q_{1},q_{2})$, meaning that, not any value of $\mu$ is allowed to generate a bistochastic matrix from the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$. The following remark clarifies how the feasible region is derived.
Remark 3.
The feasible region within the probability triangle is determined by applying a sequence of constraints arising from physical considerations. These conditions successively restrict the parameter space, resulting in a subset that satisfies all necessary properties. Figure 2 outlines how the feasible region is constructed in this case, with each subfigure depicting a specific constraint imposed on the probability triangle, as detailed below.
Figure 2(a): A subset of the probability triangle that satisfies the chain-link conditions (23) is shown in Fig. 2(a). Corresponding to each probability distribution $q=(q_{0},q_{1},q_{2})$ within the subset, the associated unitary matrix $\mathcal{U}$ generates the corresponding unistochastic matrix $\mathcal{Q}$.
Figure 2(b): For $\mu\in[1/3,1)$, the rhs of Eq. (35) can take any nonpositive value $(-\infty,0]$, but the lhs is restricted to the interval [3/2,3][-3/2,3]. This poses restrictions on the index $\mu$ in the sense that only distributions with index $\mu\in[1/3,5/9]$ can generate unistochastic matrices from $\mathcal{V}^{\dagger}\mathcal{U}$. The resulting feasible region is depicted in the Fig. 2(b).
Figure 2(c): Equations (32) to (34) pose some chain-link conditions on the four segments $q_{0},q_{1},q_{2}$, and $\sqrt{q_{i}}$ as well, i.e., the nontrivial conditions $q_{i}\leq q_{j}+q_{k}+\sqrt{q_{j}}$ hold, for $(i,j,k)$ taken from $(0,1,2)$. After imposing these constraints, the resulting feasible domain is reduced to the region shown in the Fig. 2(c).
Figure 2(d): Finally, Eq. (28), together with the condition $|\cos{\theta^{\mu}}|\leq 1$, leads to the constraint $\mu\leq 1/2$. Figure 2(d) illustrates the feasible region corresponding to $\mu\in[1/3,1/2]$, consistent with all previously imposed constraints.

### v3

@@ II.1.2 Parametrization of the distribution $q=(q_{0},\ldots,q_{d-1})$
For an arbitrary $d$-outcome probability distribution $q=(q_{0},\ldots,q_{d-1})$, we define a (d1)(d-1)-dimensional Hermitian matrix $N$ such that it has nonzero entries only along the antidiagonal, i.e., $N_{kk^{\prime}}\neq 0$ if and only if $k\oplus k^{\prime}=d$. More precisely, let $N_{kk^{\prime}}=\sum_{m=0}^{d-1}\omega^{k^{\prime}m}q_{m}=\sum_{m=0}^{d-1}\omega^{-km}q_{m}$ so that for nonzero entries we have
@@
The matrix $N$ captures the non-uniformity of the probability distribution $q=(q_{0},\ldots,q_{d-1})$ in the sense that $N=0$ if and only if the underlying distribution is uniform, i.e., $q=(1/d,\ldots,1/d)$. Clearly, $N$ has $\lceil\frac{d-1}{2}\rceil$ independent entries, where $\lceil\frac{d-1}{2}\rceil$ denotes the smallest integer greater than or equal to $\frac{d-1}{2}$, that is, it equals $\frac{d-1}{2}$ if $d$ is odd and $\frac{d}{2}$ if $d$ is even. All entries are complex, except for $N_{d/2,d/2}$ which is real and occurs when $d$ is even. With this notation, the (d1)(d-1)-dimensional matrix $N$ is fully determined by d1d-1 real independent parameters. Accordingly, by writing $N_{k,d-k}=\delta_{2k-1}+i\delta_{2k}$, we have for the (d1)(d-1) real parameters $\delta_{1},\ldots,\delta_{d-1}$
@@
for $k=1,\ldots,\lceil\frac{d-1}{2}\rceil$. This parameterization is particularly useful to quantify the extent to which the probability distribution $q=(q_{0},\ldots,q_{d-1})$ deviates from the uniform distribution $q=(1/d,\ldots,1/d)$. To illustrate the nature of this parameterization, we explicitly consider two cases of $d=2$ and $d=3$.
For $d=2$, the only nonzero entry of $N$ is given by $N_{11}=\delta$, where $\delta=q_{0}-q_{1}$ such that $1\leq\delta\leq 1$. It follows therefore that
@@
II.2 Mutually equibiased bases for $d=2$
For the simplest case of a two-level system, suppose $\mathcal{B}^{(1)}=\{|0\rangle,|1\rangle\}$ represents the standard basis. Associated with the probability distribution ${q}=(q_{0},q_{1})$, or its corresponding bistochastic matrix $\mathcal{Q}=\begin{pmatrix}q_{0}&q_{1}\\ q_{1}&q_{0}\end{pmatrix}$, we have to construct the unitary matrix $\mathcal{U}$ with prescribed moduli of its matrix elements. Starting with $\mathcal{U}=\begin{pmatrix}\sqrt{q_{0}}&\sqrt{q_{1}}\\ \sqrt{q_{1}}&\sqrt{q_{0}}{\mathrm{e}}^{i\xi}\end{pmatrix}$, unitarity requires $\xi=\pi$, hence
@@ II.2 Mutually equibiased bases for $d=2$
Although both $\mathcal{U}$ and $\mathcal{V}$ generate the same bistochastic matrix $\mathcal{Q}$ and both define bases that are MEBs with respect to the standard basis, they are not necessarily MEBs with respect to each other. For this to hold, $\mathcal{U}$ and $\mathcal{V}$ must be mutually equibiased, i.e., the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$ must generate the same bistochastic matrix $\mathcal{Q}$. This places a constraint on $\alpha$ such that $q_{0}^{2}+q_{1}^{2}+2q_{0}q_{1}\cos{\alpha}=q_{0}$, which uniquely determines $\alpha$ as
@@
with the unbiasedness parameter $\delta=q_{0}-q_{1}$ defined by Eq. (7).
Starting from the standard orthonormal basis $\{|e_{k}^{(1)}\rangle=|k\rangle\}$, one can apply the unitary matrices $\mathcal{U}$ and $\mathcal{V}$ to construct new bases $\{|e_{k}^{(2)}\rangle=\mathcal{U}|e_{k}^{(1)}\rangle\}$ and $\{|e_{k}^{(3)}\rangle=\mathcal{V}|e_{k}^{(1)}\rangle\}$ for $k=0,1$. This notion of equibiasedness admits an interesting geometric interpretation in the Bloch vector representation, in the sense that with $P_{k}^{(\alpha)}=({\mathbb{I}}_{2}+\boldsymbol{r}_{k}^{(\alpha)}\cdot\boldsymbol{\sigma})/2$ as the Bloch vector representation of the projection operator $P_{k}^{(\alpha)}$, the Bloch vectors satisfy
@@
Figure 1: Bloch vector representation for MEB, when $d=2$. Geometrically, equibiasedness means equal angle, $\vartheta=\cos^{-1}[\delta]$ between Bloch vectors of different bases.
II.3 Mutually equibiased bases for $d=3$
@@
The unitarity condition $\mathcal{U}\mathcal{U}^{\dagger}={\mathbb{I}}$, i.e., the mutual orthogonality of the rows (and columns), gives simply the conditions
@@
Equations (19)-(21) imply that the unitarity is fulfilled if we find the phases in such a way that three line segments $L_{0}$, $L_{1}$, and $L_{2}$ form a triangle. This imposes the triangle inequality on the line segments ZyczkowskiCMP2005
@@
where $(i,j,k)$ are different choices of $(0,1,2)$. For probabilities $q=(q_{0},q_{1},q_{2})$ for which the above inequalities hold, Eqs. (19)-(21) give the phases as
@@
For each probability distribution $q=(q_{0},q_{1},q_{2})$ satisfying the chain-link conditions (23), the associated unitary matrix $\mathcal{U}$, defined in Eq. (18) with phases specified in Eqs. (24)-(26), generates the corresponding unistochastic matrix $\mathcal{Q}$. By construction, the matrix $\mathcal{U}$ represents the most general form consistent with $q=(q_{0},q_{1},q_{2})$ and defines the second orthonormal basis $\{|e_{k}^{(2)}\rangle=\mathcal{U}|e_{k}^{(1)}\rangle\}$ with respect to the standard basis $\{|e_{k}^{(1)}\rangle\}$. Given $\mathcal{U}$, we have to construct two unitary matrices, $\mathcal{V}$ and $\mathcal{W}$ such that all bases are mutually equibiased. The following proposition provides a set of four MEBs for $d=3$.
@@
Starting from the standard basis $\{|e_{k}^{(1)}\rangle\}_{k=0}^{2}$, the unitary matrices $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{W}$ generate a one-parameter class of four MEBs. Here $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$ and $\mathcal{W}=\mathcal{D}_{\beta}\mathcal{U}$, where $\mathcal{D}_{\alpha}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\alpha_{1}},{\mathrm{e}}^{i\alpha_{2}}\}$ and $\mathcal{D}_{\beta}={\mathrm{diag}}\{1,{\mathrm{e}}^{i\beta_{1}},{\mathrm{e}}^{i\beta_{2}}\}$, respectively, and $\mathcal{U}$ is defined by Eq. (18). The pairs $(\alpha_{1},\alpha_{2})$ and $(\beta_{1},\beta_{2})$ are distinct and take their values from the set
@@
The proof of the proposition is given in Appendix A; here we provide a brief description of the structure of the obtained MEBs. In Fig. 2 we illustrate the one-parameter class of solutions within the feasible region. The feasible region, shown in yellow, is bounded by three hypocycloids each corresponding to the saturation of one of the chain-link inequalities (23), and three circular arcs of radius $\mu=1/2$. The hypocycloids are obtained from the condition that the matrix $\mathcal{U}$ in Eq. (18) is unitary, generating the unistochastic matrix $\mathcal{Q}$ with entries $\mathcal{Q}_{kl}=q_{k\oplus l}$. The three circular arcs, on the other hand, follow from the condition that the unitary matrix $\mathcal{V}^{\dagger}\mathcal{U}$, with $\mathcal{V}=\mathcal{D}_{\alpha}\mathcal{U}$, must generate the same unistochastic matrix $\mathcal{Q}$, up to a permutation of its entries. The latter imposes further constraints on the solutions, restricting them to a one-parameter class of probability distributions, as stated in the proposition. The class consists of six curves, each starting at $\mu=1/3$ at the center of the feasible region and ending at $\mu=1/2$ on the boundary. The curve colored in darker blue, for example, corresponds to $q=(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$ when $\alpha_{1}=0$ and $\alpha_{2}=\cos^{-1}[-\mu/(1-\mu)]$, and all other curves can be obtained by any permutation of its probability components.
@@
Figure 2: Geometrical representation of the probability distribution $q=(q_{0},q_{1},q_{2})$, for which a set of four MEBs exist. The figure is plotted in the $\delta_{2}\delta_{1}$ plane, where independent parameters $\delta_{1}$ and $\delta_{2}$ are defined in Eq. (8). The curves represent the allowed values of the distribution $q=(q^{\mu}_{+},q^{\mu}_{0},q^{\mu}_{-})$, or any permutation of its probabilities, when $\alpha_{1}=0$ and $\alpha_{2}=\cos^{-1}[-\mu/(1-\mu)]$. The circles correspond to regions with constant index $\mu$. Corresponding to each $\mu$, only six points, i.e., the distribution $q$ and its permutations, are allowed for a given $\mu$.
@@
It is worth noting that, although all three unitary matrices $\mathcal{U}$, $\mathcal{V}$, and $\mathcal{W}$ generate the same unistochastic matrix $Q$, associated with the probability distribution $q=(q_{0},q_{1},q_{2})$, the unistochastic matrices resulting from their mutual products $\mathcal{V}^{\dagger}\mathcal{U}$, $\mathcal{W}^{\dagger}\mathcal{U}$, and $\mathcal{W}^{\dagger}\mathcal{V}$ are not necessary equal to $Q$. However, they are equivalent in the sense that they are generated either by $q=(q_{0},q_{1},q_{2})$ or by a permutation of its components, such as $q^{(012)}=(q_{1},q_{2},q_{0})$ and $q^{(210)}=(q_{2},q_{1},q_{0})$, where $(012)$ and $(210)$ are forward and backward cycles of the group of permutation of three elements $\{0,1,2\}$. All such possibilities are summarized in Table 2 in Appendix A.
Note that, in Proposition 2, we assume that any two unitary matrices $\mathcal{U}$ and $\mathcal{V}$ generating the same unistochastic matrix $\mathcal{Q}$ can be written as $\mathcal{V}=\mathcal{D}_{1}\mathcal{U}\mathcal{D}_{2}$ for some diagonal unitary matrices $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$. However, for $d=3$, the existence of inequivalent unitary matrices cannot be excluded; therefore, whether this one-parameter class is exhaustive remains an open question.